OpenStudy (anonymous):

5x=2y-7 find the x and y intercept

OpenStudy (anonymous):

You can always find \(x\)-intercept by putting \(y=0\).. So, put y = 0, and find x..

OpenStudy (anonymous):

And when you again will put x = 0 and find y, that will be your y-intercept...

OpenStudy (anonymous):

5/7?

OpenStudy (anonymous):

Put y = 0, and show me what you got..

OpenStudy (anonymous):

x=5/7 y=?

OpenStudy (anonymous):

How?? I said just put y = 0, and show me, don't solve it further..

OpenStudy (anonymous):

5x=-7

OpenStudy (anonymous):

See, here you have : \(5x = 2y - 7\) Put \(y=0\), \(5x = 2(0) - 7\) \(\implies 5x = -7\) Find \(x\) now..

OpenStudy (anonymous):

If found x, then in the same way, do it again but at that time, put x = 0 and find y, that will be your y-intercept..

OpenStudy (anonymous):

x=-5/7

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

This is the basic concept dear, you are solving the next part before learning the previous part.. This is dangerous..

OpenStudy (anonymous):

\[5x = -7\] Divide by \(5\) both the sides and tell me what will you get?

OpenStudy (anonymous):

-7/5

OpenStudy (anonymous):

It is good...

OpenStudy (anonymous):

Similarly when we put : \(x=0\), \(5x = 2y - 7\) \(\implies 0 = 2y - 7\)

OpenStudy (anonymous):

7/2

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[x-intercept : (\frac{-7}{5},0) \\ y-intercept : (0, \frac{7}{2})\]